# A kinetic theory approach to consensus formation in financial markets

### v1

@@ Abstract.
We investigate the relationship between analysts’ one-year consensus forecasts for the S&P 500 index and its current level. Contrary to the conventional view that sentiment drives price forecasts, our analysis predicts that the average consensus has no effect on forecasts, while the current S&P 500 index level alone is sufficient to anticipate analysts’ price expectations.
Employing a kinetic theory framework, we model the dynamics of analysts’ opinions, by taking into account both the mutual influences shaping price consensus and the dynamics of the actual S&P 500 index level. Testing the model on real data on a long-time series of data shows that just three free parameters are enough to accurately describe the one-year average price forecasts of analysts.
@@ 1. Introduction
Opinion dynamics play an important role in understanding the emergence of collective behaviors in markets, particularly under external influences such as price movements. Conventional perspectives often link price evolution to shifts in market sentiment. In contrast, our work investigates the reverse relationship: how market sentiment reacts to and closely follows price dynamics. This approach is essential for unraveling the feedback loops that drive decision-making within financial systems.
@@
The application of methods from statistical mechanics to model social and economic systems has a long and successful history. This interdisciplinary approach was pioneered by Weidlich [15] and further popularized by Galam, Gefen, and Shapir [6]. By treating individuals as interacting entities and modeling their collective behavior, researchers have gained insights into complex social phenomena. In particular, kinetic theory has emerged as a powerful tool for describing the evolution of distributions in such systems, especially in the case of continuous opinion variables.
@@
Indeed, kinetic theory provides a natural framework for modeling the consensus formation process in an interacting community. Originally developed in statistical physics to describe the evolution of particle distributions in a gas, kinetic models have been successfully applied to various domains, including opinion dynamics and economic interactions. These models typically describe the time evolution of a probability distribution, capturing the stochastic nature of agent interactions and the emergence of macroscopic patterns from microscopic rules (see, for example, [1, 2, 3, 8, 9, 10, 14]).
@@
In the context of financial forecasting, analysts continuously update their expectations based, for example, on new information, market trends, and (possibly) peer influence. This process can be viewed as an interaction-driven system, where opinions evolve through repeated exchanges and adjustments. Unlike deterministic approaches that yield a single forecasted value, kinetic models allow for a richer description by capturing the full distribution of expectations, thus accounting for heterogeneity in analysts’ predictions. Furthermore, such models can describe key phenomena such as polarization, clustering of forecasts, and the convergence toward consensus.
@@
A kinetic approach is particularly well suited for studying the expected level of the S&P 500 index at a one-year horizon, as it allows us to incorporate both endogenous factors, such as analysts’ interactions, and exogenous influences, including macroeconomic shocks and financial news. By extending methods from kinetic theory to this setting, we aim to provide a realistic and flexible model for predicting market sentiment, ultimately improving our understanding of how collective financial expectations evolve over time.
@@
Specifically, we focus on the one-year consensus forecasts of analysts aggregated across all stocks belonging in the S&\&&P 500 index. We first demonstrate that the S&\&&P 500 and its consensus target price, more precisely, the logarithms of these two series, are cointegrated. This property is easily confirmed by using Engle-Granger procedure (see [5]). The short-run dynamics is also investigated thanks to Johansen procedure (see [11] and [12]). While it is not unexpected that these two series share a common trend, we find, surprisingly, that Granger causality test leads to the conclusion that the S&\&&P 500 level causes the consensus target price, but not vice versa. This result appears to be contradictory with several studies on the impact of analyst revisions on prices (see for instance [4], [13]). It is important to note, however, that these authors focus on analyst earnings revisions, while we examine their ability to forecast prices at a one-year horizon. Furthermore, these studies focus on individual stocks, while we are interested in the aggregate forecast for the entire S&\&&P 500.
@@
Consequently, a central hypothesis of our study, while designing the kinetic model, is that the consensus does not predict or influence market movements, but rather follows them. Our model assumes that two main factors drive the analysts’ opinions: the influence of other analysts (described as a mean-field interaction) and the actual price of the index. This framework allows us to explore the interplay between these two effects and their impact on opinion formation.
@@
Furthermore, our analysis shows that the influence of the mean-field becomes negligible over time, leaving the current price as the sole determinant of the average one-year forecasts. Hence, the average of the analysts’ one-year forecasts on the S&P 500 index can be explained using only three fixed parameters: $\Delta$, representing a percentage of the current price; $q$, a coefficient for the convex combination of the current index value and the analyst’s current opinion; and $\beta$, the relative weight of following the index value versus the mean-field.
@@
2. exploratory data analysis
2.1. Datas
Our study is based on the one-year target price published by Bloomberg for the S&P 500 across all analysts. More specifically, bank or broker analysts publish their one-year target prices for equities, particularly those in the S&P 500. Bloomberg collects and distributes these datas to market participants via its terminal. Then, it calculates and publishes a one-year consensus forecast for each stock based on the median of the whole one-year target prices. Bloomberg finally computes the sum of one-year consensus for S&P 500 stocks weighted by the weight of each stock in the index to obtain a one-year consensus forecast for the level of the S&P 500 based on analyst recommendations. The daily data analyzed in this article cover the period from May 26th, 2009, to November 20th, 2024.
Figure 1. Time series of the S&P 500 index $X(t)$ and the one-year analysts’ forecasts for the S&P 500 index, from May 26th, 2009, to November 20th, 2024. Source: Bloomberg
In the remainder of this paper, $X(t)$ will denote the level of the S&P 500 at time $t$ while $s(t)$ will be the one year consensus target price computed by Bloomberg for the S&P 500.
2.2. Long-run relationship between log price and log target price for S&P 500
It is quite obvious that $\ln(X(t))$ and $\ln(s(t))$ are nonstationary and, indeed, stationarity is rejected by the Augmented Dickey-Fuller test at any confidence level.
For a discrete time series $Y(t)$, define the first difference $DY(t)=Y(t)-Y(t-1)$. This time, $D\ln(X(t))$ and $D\ln(s(t))$ are accepted to be stationary according to the ADF test (at $1\%$ confidence level). The time series $\ln(X(t))$ and $\ln(s(t))$ are then said to be $I(1)$ (integrated of order 1) and their first differences, which are stationary time series, are said to be $I(0)$ (integrated of order 0).
Now, we investigate the cointegration between $\ln(X(t))$ and $\ln(s(t))$, which is a linear relation linking theses series. In that particular case, the cointegration relationship would be defined as follows:
@@ 2.2. Long-run relationship between log price and log target price for S&P 500
Usually, one cannot use the Ordinary Least Square method to estimate parameters in the cointegration relationship (1) for it leads to false conclusions. This issue has first been exhibited by Yule (see [16]). Decades later, Granger and Newbold investigated this misprocedure more deeply [7] and baptised it “spurious regression”. Thereafter, Engle &\&& Granger ([5]) proposed the following procedure to investigate cointegration and estimate the parameters whenever $Y(t)$ and $Z(t)$ are proved to be $I(1)$:
@@
The estimation of the linear regression $\ln(s(t))=\alpha\ln(X(t))+\beta+u(t)$ leads to $\widehat{\alpha}=0.98679$ and $\widehat{\beta}=0.21778$.
Augmented Dickey-Fuller test allows us to accept stationarity for $e(t)=\ln(s(t))-0.98679\ln(X(t))-0.21778$ (at $1\%$ confidence level).
This leads to the conclusion that $\ln(X(t))$ and $\ln(s(t))$ are cointegrated. This appears clearly by looking at figure 2 which compares $\ln(s(t))$ with its estimated long-term equilibrium level $\alpha\ln(X(t))+\beta$.
Figure 2. Logarithm of one-year analysts’ forecasts for the S&P 500 index vs its long-term equilibrium level, from May 26th, 2009, to November 20th, 2024.
@@ 2.3. Short-run relationship
One major result regarding cointegration theory is the Granger-Engle Representation Theorem (see [5]):
@@
The name of Model (2) stems from the fact that the error in the cointegration relationship at time $t-1$ is corrected at time $t$ by the restoring force $\gamma$, which must be negative. The more the “error” at time $t-1$ in the cointegration relationship is large, the stronger the restoring force is.
@@
Hence, we have to investigate two short-term equations, but we keep the same long-run relationship in both models, $s(t)$ being the response variable. Consequently, $\gamma^{\prime}$ estimated below will remain negative while $\gamma$ will become positive. Based on three out of four information criteria (Akaike, Schwarz, Hannan-Quin and Final Prediction error), the optimal number of lags $p$ for short-run dynamics is estimated to be 9 (see Table 1). This is therefore the value that will be retained to estimate the short-term equations.
@@
Table 1. Optimal number of lags for each criterion
@@ 2.3.1. Short-run relationship with $D\ln(X(t))$ as the dependent variable
The first model we look at is:
$D\ln\left(X(t)\right)=\sum_{i=1}^{9}a_{i}D\ln(X(t-i))+\sum_{j=1}^{9}b_{j}D\ln(s(t-j))+\gamma\left(\ln(s(t-1))-\alpha\ln(X(t-1))-\beta\right)+\varepsilon(t).$
The model is estimated in R according to Johansen procedure by using cajorls and vecm functions. None of the coefficients of the past level of $D\ln(s(t))$ (i.e. from $b_{1}$ to $b_{9}$) are significant (see Table 2 below). Hence $\ln(s(t))$ does not (Granger-)cause $\ln(X(t))$. The $R^{2}$ is only $2\%$, which is very small. We can deduce that the short-term dynamics of the market is not impacted at all by the consensus. One can remark, though, that there is a bit of autocorrelation for $\ln(X(t))$ (e.g. $a_{1}$ is significant).
@@
$b_{8}$0.266791 0.219320 -1.216
$b_{9}$0.138858 0.205359 -0.676
***, **, *, $\cdot$ indicate stastistical significance at $0.1\%$, $1\%$, $5\%$, $10\%$ levels respectively
2.3.2. Short-run relationship with $D\ln(s(t))$ as the dependent variable
The second model is written:
$D\ln(s(t))=\sum_{i=1}^{9}a^{\prime}_{i}D\ln(s(t-i))+\sum_{j=1}^{9}b^{\prime}_{j}D\ln(X(t-j))+\gamma^{\prime}\left(\ln(s(t-1))-\alpha\ln(X(t-1))-\beta\right)+\varepsilon^{\prime}(t).$
The short-term equation with $D\ln(s(t))$ as the dependent variable is reported in Table 3. The $R^{2}$ of the model is around $60\%$, wich is quite good, and most of the coefficients are significant.

### v2

@@ Abstract.
It is sometimes acknowledged that (sell-side) equity analysts’ recommendations influence investors and therefore market prices. In particular, the S&P 500 is expected to decline (respectively rise) when analysts revise their targets downward (respectively upward). Our findings indicate not only that analysts’ consensus exert no influence on market prices, but also that, conversely, analysts appear to set their target prices based on markets prices.
Employing a kinetic theory framework, we model the dynamics of analysts’ opinions, by taking into account both the mutual influences shaping price consensus and the dynamics of the actual S&P 500 index level. The model is calibrated on a training subset of data and tested on an independent set to assess its predictive power. Our tests show that just three free parameters are enough to accurately predict the one-year average price forecasts of analysts.
Key words and phrases:
S&P 500; Cointegration; Kinetic modeling; Forecast analysis. JEL Codes: C63
@@ 1. Introduction
Opinion dynamics play an important role in understanding the emergence of collective behaviors in markets, particularly under external influences such as price movements. Conventional perspectives often link price evolution to shifts in analysts’ recommendations. In contrast, our work investigates the reverse relationship: how analysts’ recommendations react to and closely follows price dynamics. This approach is essential for unraveling the feedback loops that drive decision-making within financial systems.
@@
The application of methods from statistical mechanics to model social and economic systems has a long and successful history. This interdisciplinary approach was pioneered by Weidlich [20] and further popularized by Galam, Gefen, and Shapir [9]. By treating individuals as interacting entities and modeling their collective behavior, researchers have gained insights into complex social phenomena.
@@
Indeed, kinetic theory, originally developed in statistical physics to describe the evolution of particle distributions in a gas, provides a natural framework for modeling the consensus formation process in an interacting community [17], [2].
Kinetic models typically describe the time evolution of a probability distribution, capturing the stochastic nature of agent interactions and the emergence of macroscopic patterns from microscopic rules (see, for example, [1], [3], [4], [5], [11], [12], [13], [19]).
This paper investigates consensus formation within the field of financial analysis. Employed by financial institutions such as investment banks or brokerage firms, sell-side equity analysts are tasked with evaluating the financial performance of publicly traded companies. Their primary objective is to formulate investment recommendations for equity investors. These recommendations typically take the form of ’buy,’ ’hold’ (or ’neutral’), or ’sell’ ratings, often accompanied by a one-year price target for the covered securities. It is crucial to note that equity analysts are not investors themselves; indeed, they are explicitly prohibited from trading in the shares of the companies they cover. Their market influence, therefore, is derived exclusively from the market’s reaction to the dissemination of their recommendations, given that a substantial portion of investors act upon these advisories.
@@
In the context of financial forecasting, analysts continuously update their expectations based, for example, on new information, market trends, and (possibly) peer influence. This process can be viewed as an interaction-driven system, where opinions evolve through repeated exchanges and adjustments. Unlike deterministic approaches that produce a single predicted value, kinetic models allow for a richer description by capturing the full distribution of expectations, thus accounting for heterogeneity in analysts’ predictions.
For instance, it allows us to compute higher-order moments such as the variance, as detailed in Section 4, which are crucial for understanding uncertainty and predicting risk or volatility.
Furthermore, such models can describe key phenomena such as polarization, clustering of forecasts, and the convergence toward consensus.
@@
A kinetic approach is particularly well suited for studying the expected level of the S&P 500 index at a one-year horizon, as it allows us to incorporate both endogenous factors, such as analysts’ interactions, and exogenous influences, including macroeconomic shocks and financial news. By extending methods from kinetic theory to this setting, we aim to provide a realistic and flexible model for predicting analysts’ recommendations, ultimately improving our understanding of how collective financial expectations evolve over time.
@@
Specifically, we focus on the one-year consensus forecasts of analysts aggregated across all stocks belonging in the S&\&P 500 index. We first demonstrate that the S&\&P 500 and its consensus target price, more precisely, the logarithms of these two series, are cointegrated. This property is easily confirmed by using Engle-Granger procedure [8]. The short-run dynamics is also investigated thanks to the
Johansen-Juselius multivariate cointegration technique [14], [15] and [16].
While it is not unexpected that these two series share a common trend, we find, surprisingly, that Granger causality test leads to the conclusion that the S&\&P 500 level causes the consensus target price, but not vice versa. This result appears to be contradictory with several studies on the impact of analyst revisions on prices (see for instance [6], [18]). It is important to note, however, that these authors focus on analyst earnings revisions, while we examine their ability to forecast prices at a one-year horizon. Furthermore, these studies focus on individual stocks, while we are interested in the aggregate forecast for the entire S&\&P 500.
@@
Consequently, a central hypothesis of our study, while designing the kinetic model, is that the analysts’ consensus does not predict or influence market movements, but rather follows them. Our model assumes that two main factors drive the analysts’ opinions: the influence of other analysts (described as a mean-field interaction) and the actual price of the index. This framework allows us to explore the interplay between these two effects and their impact on opinion formation.
@@
Furthermore, our analysis shows that the influence of the mean-field becomes negligible over time, leaving the current price as the sole determinant of the average one-year forecasts. Hence, the average of the analysts’ one-year forecasts on the S&P 500 index can be modeled using only three fixed parameters: $\Delta$, representing the relative premium applied as a proportional adjustment to the current index level; $q$, a convex combination coefficient weighting the current index value $X(t)$, augmented by a spread, and the analyst’s prior forecast; and $\beta$, modulating the relative influence of the index dynamics with respect to the mean-field consensus.
To validate the robustness of our approach, we divided the dataset into two parts: a training set, used to estimate the model’s parameters, and a testing set, used to evaluate its predictive performance on data not seen during calibration.
@@
2. Exploratory data analysis based on Engle and Granger approach
2.1. Data
Our study uses the one-year target prices for S&\&P 500 stocks published by Bloomberg. These targets are provided by equity analysts at banks and brokerages.
Bloomberg aggregates this data and distributes it to market participants through its terminal. The company then calculates a one-year consensus forecast for each stock, using the average of all individual target prices. Bloomberg finally computes the sum of one-year consensus for S&P 500 stocks weighted by the weight of each stock in the index to obtain a one-year consensus forecast for the level of the S&P 500 based on analysts’ recommendations. Equity analysts are typically required to publish reports on a regular basis, at least once per quarter. Nevertheless, they also provide commentary whenever events are expected to influence the price of the stocks they cover. As a result, an analyst may revise his target price several times within a year. More than 30 banks and brokerage firms contribute to Bloomberg, generally the biggest ones. Given that the S&\&P 500 is composed of 500 stocks, the aggregate target price is subject to daily fluctuations, although the magnitude of these changes may be small. The daily data analyzed in this article cover the period from May 26th, 2009, to June 2nd, 2025.
Figure 1. Time series of the S&P 500 index $X(t)$ (dark grey line) and the one-year analysts’ forecasts for the S&P 500 index (grey line), from May 26th, 2009, to June 2nd, 2025. Source: Bloomberg
In the remainder of this paper, $X(t)$ will denote the level of the S&P 500 at time $t$ while $s_{\text{meas}}(t)$ will be the one year consensus target price computed by Bloomberg for the S&P 500.
2.2. Test for stationarity
It is quite obvious that $\ln(X(t))$ and $\ln(s_{\text{meas}}(t))$ are nonstationary and, indeed, stationarity is rejected by the Augmented Dickey-Fuller test at any confidence level (cf table 1).
Table 1. Augmented Dickey-Fuller test for $\ln(X(t))$ and $\ln(s_{\text{meas}}(t))$
t-statistic Prob
ADF test statistics for $\ln(X(t))$ 0.74 0.34
ADF test statistics for $\ln(s_{\text{meas}}(t))$ 2.35 0.62
Test critical values 1% level -2.58
Test critical values 5% level -1.95
Test critical values 10% level -1.62
For a discrete time series $Y(t)$, define the first difference DY(t)=Y(t)Y(t1)DY(t)=Y(t)-Y(t-1). This time, $D\ln(X(t))$ and $D\ln(s_{\text{meas}}(t))$ are accepted to be stationary according to the ADF test (cf table 2).
Table 2. Augmented Dickey-Fuller test for $D\ln(X(t))$ and $\ln(s_{\text{meas}}(t))$
t-statistic Prob
ADF test statistics for $D\ln(X(t))$ -8.26 <<0.01
ADF test statistics for $D\ln(s_{\text{meas}}(t))$ -4.62 << 0.01
Test critical values 1% level -2.58
Test critical values 5% level -1.95
Test critical values 10% level -1.62
Finally, the time series $\ln(X(t))$ and $\ln(s_{\text{meas}}(t)$ are accepted as $I(1)$ (integrated of order 1) because they are non-stationary series whose first differences are stationary (or $I(0)$).
2.3. Cointegration test
Now, we investigate the cointegration between the quantities $\ln(X(t))$ and $\ln(s_{\text{meas}}(t))$, which is a linear relation linking theses series. In that particular case, the cointegration relationship would be defined as follows:
@@ 2.3. Cointegration test
Usually, one cannot use the Ordinary Least Square method to estimate parameters in the cointegration relationship (1) for it leads to false conclusions. This issue has first been exhibited by Yule (see [21]). Decades later, Granger and Newbold investigated this misprocedure more deeply [10] and baptised it “spurious regression”. Thereafter, Engle &\& Granger ([8]) proposed the following procedure to investigate cointegration and estimate the parameters whenever $Y(t)$ and $Z(t)$ are proved to be $I(1)$:
@@
The estimation of the linear regression $\ln(s_{\text{meas}}(t))=\alpha\ln(X(t))+\beta+u(t)$ leads to $\widehat{\alpha}=0.98679$ and $\widehat{\beta}=0.21778$. Augmented Dickey-Fuller test allows us to accept stationarity for $e(t)=\ln(s_{\text{meas}}(t))-0.98679\ln(X(t))-0.21778$ (cf table 3).
Table 3. Augmented Dickey-Fuller test applied to the residuals from the regression $\ln(s_{\text{meas}}(t))$ on $\ln(X(t))$
t-statistic Prob
ADF test statistics for $e(t)$ -2.82 <<0.01
Test critical values 1% level -2.58
Test critical values 5% level -1.95
Test critical values 10% level -1.62
This leads to the conclusion that $\ln(X(t))$ and $\ln(s_{\text{meas}}(t))$ are cointegrated. This appears clearly by looking at figure 2 which compares $\ln(s_{\text{meas}}(t))$ with its estimated long-term equilibrium level $\alpha\ln(X(t))+\beta$.
Figure 2. Logarithm of one-year analysts’ forecasts for the S&P 500 index (dark grey line) vs its long-term equilibrium level (grey line), from May 26th, 2009, to June 2nd, 2025.
@@ 2.4. Short-run relationship
One major result regarding cointegration theory is the Granger-Engle Representation Theorem [8]:
@@
The name of Model (2) stems from the fact that the error in the cointegration relationship at time t1t-1 is corrected at time $t$ by the restoring force $\gamma$, which must be negative. The more the “error” at time t1t-1 in the cointegration relationship is large, the stronger the restoring force is.
@@
Accordingly, we need to estimate two short-term equations simultaneously. To investigate this further, we will use the Johansen procedure, as detailed in the following section.
Remark 2.5.
The long-run relationship in the same in both short-term models, $s_{\text{meas}}(t)$ being the response variable. Consequently, $\gamma^{\prime}$ estimated below will remain negative while $\gamma$ will become positive.
3. Johansen procedure
We use the Johansen procedure with Maximum Likelihood Estimation (MLE) to determine the parameters of the Vector Error Correction Model (VECM). This estimation is performed in the R programming environment using cajorls and vecm functions. A crucial initial step is testing the rank, $r$, of the VECM’s long-run coefficient matrix $\Pi$. This rank $r$ represents the number of distinct cointegration relationships. The Johansen procedure is preferred over the Engle-Granger approach because it can simultaneously handle multiple variables and relationships. However, since our analysis is limited to only two variables, ($\ln(s_{\text{meas}}(t))$ and $\ln(X(t))$), we can have at most one cointegration relationship $(r\leq 1)$.
3.1. Tests for lags and cointegration
Based on three out of four information criteria (Akaike, Schwarz, Hannan-Quin and Final Prediction error), the optimal number of lags $p$ for short-run dynamics is estimated to be 9 (see Table 4). This is therefore the value that will be retained to estimate the model.
Table 4. Optimal number of lags for each criterion
@@ 3.1. Tests for lags and cointegration
The need for a large number of lags stems from the large residual outliers observed in some sub-periods of the cointegration system. This is a recognized characteristic of equity markets: their frequent experience of volatility regime changes. However, we confirmed that the optimal number of lags decreases with the sample size; for instance, restricting the analysis to the first three years of historical data only requires four lags, as reported in Table 5.
Table 5. Optimal number of lags for each criterion (2009-2012)
AIC SIC HQ FPE
4 3 4 4
If the estimation is restricted to the first year of the historical data, the necessary number of lags is lowered even further, as reported in Table 6.”
Table 6. Optimal number of lags for each criterion (2009-2010)
AIC SIC HQ FPE
2 2 2 2
To test the number of cointegration relationship, there are two methods (trace and maximum eigenvalue), but again, these are identical if we only deal with two variables. Now, the cointegration rank test lead to the result reported in table (7).
Table 7. Johansen rank test
test 10% 5% 1%
$r\leq 1$ 11.81 7.52 9.24 12.97
$r=0$ 178.74 13.75 15.67 20.20
[…]
3.2. Estimates
Even though the Johansen procedure allows for the estimation of both the short-run and the long-run relationships in one step, we are going to comment on each relationship individually.
3.2.1. Long-term relationship
The estimation leads to:
$\ln(s_{\text{meas}}(t))-0.99186\ln(X(t))-0.20574\leadsto I(0).$
This result is very close to the estimate obtained earlier using the simpler Engle-Granger approach. It is this term that provides the mechanism to correct for errors in the short-run model.
@@ 3.2.2. Short-run relationship with $D\ln(X(t))$ as the dependent variable
The equation is the following:
$D\ln\left(X(t)\right)=\sum_{i=1}^{9}a_{i}D\ln(X(t-i))+\sum_{j=1}^{9}b_{j}D\ln(s_{\text{meas}}(t-j))+\gamma\left(\ln(s_{\text{meas}}(t-1))-\alpha\ln(X(t-1))-\beta\right)+\varepsilon(t).$
None of the coefficients of the past level of $D\ln(s_{\text{meas}}(t))$ (i.e. from $b_{1}$ to $b_{9}$) are significant (see Table 8 below). Hence $\ln(s_{\text{meas}}(t))$ does not (Granger-)cause $\ln(X(t))$. The $R^{2}$ is only $2\%$, which is very small. We can deduce that the short-term dynamics of the market is not impacted at all by the consensus price. One can remark, though, that there is a bit of autocorrelation for $\ln(X(t))$ (e.g. $a_{1}$ is significant).
***, **, *, $\cdot$ indicate statistical significance at $0.1\%$, $1\%$, $5\%$, $10\%$ levels respectively
@@
$b_{8}$ -0.266791 0.219320 -1.216
$b_{9}$ -0.138858 0.205359 -0.676
3.2.3. Short-run relationship with $D\ln(s_{\text{meas}}(t))$ as the dependent variable
The studied equation is the following:
$D\ln(s_{\text{meas}}(t))=\sum_{i=1}^{9}a^{\prime}_{i}D\ln(s_{\text{meas}}(t-i))+\sum_{j=1}^{9}b^{\prime}_{j}D\ln(X(t-j))+\gamma^{\prime}\left(\ln(s_{\text{meas}}(t-1))-\alpha\ln(X(t-1))-\beta\right)+\varepsilon^{\prime}(t).$
The short-term equation with $D\ln(s_{\text{meas}}(t))$ as the dependent variable is reported in Table 9. The $R^{2}$ of the model is around $60\%$, wich is quite good, and most of the coefficients are significant.
Table 9. Estimates of the short-term equation for the response variable $D\ln(s_{\text{meas}}(t))$
***, **, *, $\cdot$ indicate statistical significance at $0.1\%$, $1\%$, $5\%$, $10\%$ levels respectively